# Majorana edge and end states in planar Josephson junctions

### v2

@@ IV Summary
We investigated the formation and properties of edge-like and end-like MSs in proximitized planar JJs subjected to an in-plane magnetic field, considering the effects of Rashba and/or Dresselhaus SOCs. The end-like MSs are primarily localized at opposite ends of the normal region within the junction. In contrast, the edge-like MSs extend along the system’s edges, perpendicular to the junction. To characterize the nature and protection of the MSs we introduced a quantity called the topological gap character. This quantity provides insight into whether the system is in the D-class (or BDI-class with $\mathbb{Z}_{2}$ index equal to ±$\pm 1$) TS state, the size of the topological gap, and whether the MSs are end-like or edge-like. Through numerical simulations of the topological gap character as a function of the magnetic field strength and the superconducting phase difference ($\phi$) across the junction, we found that when the system is in the TS state, and either Rashba or Dresselhaus SOC dominates, edge-like and end-like MSs, protected by a sizable topological gap, typically emerge near $\phi$ values of 0 and $\pi$, respectively. However, when the Rashba and Dresselhaus SOC strengths are comparable, and the junction and magnetic field are properly oriented, protected edge-like MSs can also emerge at phases near $\pi$. Our results reveal the possibility of inducing topologically protected transitions between edge-like and end-like MSs in phase-biased JJs by tuning the superconducting phase difference and the magnetic field strength. Furthermore, in phase-unbiased junctions with comparable Rashba and Dresselhaus SOC strengths, protected transitions between end and edge-like MSs can be achieved by solely adjusting the magnetic field strength, as the superconducting phase self-tunes to minimize the system’s free energy. Controlled transitions between edge-like and end-like MSs could function as switchable global interconnects, enabling or disabling the coupling between distant MSs.
@@ Acknowledgments
parameter Al/HgTe Scharf et al. (2019b) Al/InSb Mayer et al. (2020)
@@
Table 1: Material parameters used in the numerical simulations of proximitized planars JJs reported in this work. The proximity-induced superconducting gap is represented by $\Delta_{0}$, $\mu_{S}$ ($\mu_{N}$) denotes the chemical potential in the S (N) region, $m^{\ast}$ is the electron effective mass (with $m_{0}$ as the bare mass of the electron), $\lambda=\sqrt{\alpha^{2}+\beta^{2}}$ characterizes the combined strength of Rashba ($\alpha$) and Dresselhaus ($\beta$) SOCs, $W_{N}$ ($W_{S}$) is the width of the N (S) region, $L$ is the length of the junction, $a$ is the TB lattice constant, and $t$ the TB hopping parameter.
@@
Figure 8: Probability density (normalized to its maximum value) for JJs with (a) $W_{N}=0.4\;\mu$m and (b) $W_{N}=0.8\;\mu$m. The other parameters were taken as in Fig. 7(c), namely, $W_{S}=1.2\;\mu$m, $E_{Z}=0.25$ meV, $\theta_{so}=\pi/4$, $\theta_{c}=3\pi/4$, $\phi=0$, and $\varphi_{B}=\pi/2$.
@@ B.1 Magnetic field along the junction ($\mathbf{E}_{Z}\parallel\hat{\mathbf{y}}$)
where $j=e=1$ ($j=h=-1$) for electron-like (hole-like) states, $E_{Z}=|\mathbf{E}_{Z}|$, and
@@ B.2 Magnetic field perpendicular to the junction ($\mathbf{E}_{Z}\perp\hat{\mathbf{y}}$)
In Sec. III we focused our discussion on narrow junctions with $W_{N}=0.1\;\mu$m [Figs. 3-6] and $W_{N}=0.2\;\mu$m. However, the existence of edge-like MSs is also supported in wider junctions as evidenced by Fig. 8, where the probability density (normalized to its maximum value) is shown for (a) $W_{N}=0.4\;\mu$m and (b) $W_{N}=0.8\;\mu$m. The other system parameters were taken as in Fig. 7(c).

### v3

@@ IV Summary
We investigated the formation and properties of edge-like and end-like MSs in proximitized planar JJs subjected to an in-plane magnetic field, considering the effects of Rashba and/or Dresselhaus SOCs. The end-like MSs are primarily localized at opposite ends of the normal region within the junction. In contrast, the edge-like MSs extend along the system’s edges, perpendicular to the junction. To characterize the nature and protection of the MSs we introduced a quantity called the topological gap character. This quantity provides insight into whether the system is in the D-class (or BDI-class with $\mathbb{Z}_{2}$ index equal to $\pm 1$) TS state, the size of the topological gap, and whether the MSs are end-like or edge-like. Through numerical simulations of the topological gap character as a function of the magnetic field strength and the superconducting phase difference ($\phi$) across the junction, we found that when the system is in the TS state, and either Rashba or Dresselhaus SOC dominates, edge-like and end-like MSs, protected by a sizable topological gap, typically emerge near $\phi$ values of 0 and $\pi$, respectively. However, when the Rashba and Dresselhaus SOC strengths are comparable, and the junction and magnetic field are properly oriented, protected edge-like MSs can also emerge at phases near $\pi$. Our results reveal the possibility of inducing topologically protected transitions between edge-like and end-like MSs in phase-biased JJs by tuning the superconducting phase difference and the magnetic field strength. Furthermore, in phase-unbiased junctions with comparable Rashba and Dresselhaus SOC strengths, protected transitions between end and edge-like MSs can be achieved by solely adjusting the magnetic field strength, as the superconducting phase self-tunes to minimize the system’s free energy. Controlled transitions between edge-like and end-like MSs could function as switchable global interconnects, enabling or disabling the coupling between distant MSs. However, this approach requires maintaining low levels of disorder, since increasing electrostatic disorder can shrink the topological gap and introduce uncertainty in the MS localization.
@@ Acknowledgments
parameter Al/HgTe Scharf2019:PRB Al/InSb mayer2020gate
@@
Table 1: Material parameters used in the numerical simulations of proximitized planars $J$s reported in this work. The proximity-induced superconducting gap is represented by $\Delta_{0}$, $\mu_{S}$ ($\mu_{N}$) denotes the chemical potential in the S (N) region, $m^{\ast}$ is the electron effective mass (with $m_{0}$ as the bare mass of the electron), $\lambda=\sqrt{\alpha^{2}+\beta^{2}}$ characterizes the combined strength of Rashba ($\alpha$) and Dresselhaus ($\beta$) SOCs, $W_{N}$ ($W_{S}$) is the width of the N (S) region, $L$ is the length of the junction, $a$ is the TB lattice constant, and $t$ the TB hopping parameter.
@@
[…]
@@ B.1 Magnetic field along the junction ($\mathbf{E}_{Z}\parallel\hat{\mathbf{y}}$)
where $j=e=1$ (j=h=1j=h=-1) for electron-like (hole-like) states, $E_{Z}=|\mathbf{E}_{Z}|$, and
@@ B.2 Magnetic field perpendicular to the junction ($\mathbf{E}_{Z}\perp\hat{\mathbf{y}}$)
[…]